A 7 m diameter spherical ball is known to emit radiation at a rate of 340 W when its surface temperature is 880 K. Determine the average emissivity of the ball at this temperature. 𝜎 = 0.1714 x10-8 W/m2K 4 .


1
Expert's answer
2021-07-22T17:47:01-0400
"e=\frac{P}{\sigma AT^4}\\
e=\frac{340}{(0.25\cdot0.1714\cdot10^{-8}) (7)^2(880)^4}=0.027"